# Qudit inspired optimization for graph coloring

### v1

@@ Abstract
We introduce a quantum-inspired algorithm for Graph Coloring Problems (GCPs) that utilizes qudits in a product state, with each qudit representing a node in the graph and parameterized by $d$-dimensional spherical coordinates. We propose and benchmark two optimization strategies: qudit gradient descent (QdGD), initiating qudits in random states and employing gradient descent to minimize a cost function, and qudit local quantum annealing (QdLQA), which adapts the local quantum annealing method to optimize an adiabatic transition from a tractable initial function to a problem-specific cost function. Our approaches are benchmarked against established solutions for standard GCPs, showing that our methods not only rival but frequently surpass the performance of recent state-of-the-art algorithms in terms of solution quality and computational efficiency. The adaptability of our algorithm and its high-quality solutions, achieved with minimal computational resources, point to an advancement in the field of quantum-inspired optimization, with potential applications extending to a broad spectrum of optimization problems.
@@ I Introduction
Today, many prominent tasks in industry can be formulated as integer optimization problems, such as how to efficiently use resources kallrath2005solving or do tasks like job scheduling ku2016mixed or portfolio optimization he2014two . One important example of integer optimization is the graph coloring problem (GCP): given a graph with $V$ nodes and $\mathcal{E}$ edges, assign colors to the nodes such that no two connected nodes share the same color. More generally, one might try to find the minimum number of colors needed for such an assignment to be possible, henceforth referred to as the chromatic number of a graph  lewis_16 . Being NP-Complete garey1974some , the GCP can be related to a large number of real-world problems, including air traffic control barnier_14 , charging electric vehicles deller_23 , resource allocation (e.g. supply-chain or portfolio management), even tasks like designing timetables, seating plans, sports tournaments or sudoku lewis_16 . Thus, there is much active research in developing algorithms that efficiently find good or optimal solutions to the GCP and scale favourably with the size of the graph.
@@
Quantum technologies might provide one avenue for efficiently solving optimization problems through algorithms like the Quantum Approximate Optimization Algorithm (QAOA) farhi_14 ; hadfield_19 ; oh_19 or quantum annealing hauke_2020 ; pokharel_21 ; yarkoni_2022 . While there is a straight-forward mapping from binary optimization problems to qubit systems lucas_14 ; glover_22 , some integer optimization problems, like GCPs, are more naturally formulated using larger local degrees of freedom. One way to accomplish this in a physics-inspired manner is to extend the notion of a qubit to a $d$-dimensional local degree of freedom, known as a qudit. This, together with the improved capacity for simulations of physical systems with a larger number of local physical degrees of freedom, has inspired the development of hardware implementing qudit systems and qudit-inspired algorithms lu_20 ; wang_20 ; ringbauer_22 . For instance, in Ref. deller_23 , Deller et al. extend QAOA and integer optimization problems, including the GCP, to the qudit regime, and Ref. bottrill_23 explores using qutrits (3-dimensional local degrees of freedom) for the GCP.
@@
Complementary to the improvement of quantum technologies, the connection between quantum physics and optimization problems has enhanced the interest in so-called “physics-inspired algorithms” for optimization. A variety of such algorithms have emerged based on techniques like mean field theory smolin_14 ; veszeli_22 ; bowles_22 , tensor networks bauer_15 ; mugel_22 , neural networks gomes_19 ; Zhao_2021 ; schuetz_22 ; khandoker_2023 , and dynamical evolution Tiunov_19 ; goto_19 ; goto_21 . Some of these new approaches have found significant success, providing state-of-the-art solutions to existing problems as well as benchmarks to compare against possible quantum-hardware-based solutions.
@@
In this work, we introduce a qudit-inspired algorithm for solving integer optimization problems. First, our approach maps the problem to a qudit system where each qudit is parametrized by $d-$dimensional spherical coordinates. Then, we propose two schemes to solve the optimization problem. In the first scheme, which we refer to as qudit gradient descent (QdGD), each qudit is initialized to a random state and we minimize a cost function via gradient descent. Our second scheme, which we refer to as qudit local quantum annealing (QdLQA), we simulate the quantum annealing from the ground state of a simple Hamiltonian to that of a problem-specific cost function. This is the concept of local quantum annealing introduced by Bowles et al. in Ref. bowles_22 , in the sense that the system is constrained to remain in a product state through its evolution. We extend this idea to integer optimization, without requiring that the cost function at the end of the evolution has to have an interpretation as a quantum physical Hamiltonian. Our intuition for this approach is that we find the minima of a complicated function $f_{\textrm{final}}(\theta)$ by first initializing the parameters, $\theta$, to a known minimum of a simple function $f_{\textrm{initial}}(\theta)$. Then, by slowly changing $f_{\textrm{initial}}(\theta)\rightarrow f_{\textrm{final}}(\theta)$ while minimizing with respect to the parameters $\theta$, an optimal or good solution to $f_{\textrm{final}}(\theta)$ is found.
@@
We test the algorithms on different standard GCPs and compare to different benchmarks in the literature hebrard_19 ; schuetz_22_2 ; wang_23 . In almost all cases, our approach obtains results better than or equal to recently proposed state-of-the-art algorithms such as graph-neural-network-based approaches schuetz_22_2 ; wang_23 . We also run our algorithms on large sparse datasets where many colors/local degrees of freedom (up to 70 hebrard_19 ) are required to solve the problem. We find that even though our approaches do not find the global minimum for such high-dimensional problems, they still provide relatively good solutions (errors of the order $10^{-3}$) for problem sizes unaccounted for in state-of-the-art works such as schuetz_22_2 ; wang_23 . Notably, in the case of QdLQA, we find that all of these results can be obtained doing only a few (for most problems only one) gradient descent steps at each point in the adiabatic evolution, as was achieved in Ref. bowles_22 .
@@
The rest of this document is structured as follows: In Sec. II, we introduce the GCP and the algorithms. In Sec. III we benchmark our approach with other state-of-the-art methods, and in Sec. IV, we provide a summary and a discussion of improvements and further applications.
@@ II.1 Graph coloring
In our approach, we aim to solve the GCP by re-framing it as finding the ground state of the antiferromagnetic Potts model potts_1952 ; wu_82 . The two are closely related wu_88 ; Zdeborova_07 ; schuetz_22_2 , and for an undirected graph $G=(V,\mathcal{E})$ with a set of nodes $V$, edges $\mathcal{E}$, and a set of $c$ colors $\{0,...,c-1\}$, the energy is given by
@@
where each spin is set to one of the colors $\sigma_{i}\in\{0,...,c-1\}$ and $\delta(\sigma_{i},\sigma_{j})=1$ if $\sigma_{i}=\sigma_{j}$ and $\delta(\sigma_{i},\sigma_{j})=0$ otherwise. Thus, this energy is minimized whenever as many possible spins have a different color to each of their neighbours (thus as many terms as possible being zero). The chromatic number, $\chi(G)$, is defined as the minimum number of colors $c$ required for this cost to reach zero, i.e., when no neighbouring vertices share the same color. This problem can be posed in a number of forms, either to simply minimize the cost (NP-complete), to check whether a given graph admits a coloring with a certain number of colors (NP-complete), or to find the chromatic number of a graph (NP-hard) garey1974some .
@@ II.2 Qudit ansatz
|$\ket{\psi}_{i}=\vec{\psi}_{i}=\begin{bmatrix}\cos(\phi^{i}_{0})\\ \sin(\phi^{i}_{0})\cos(\phi^{i}_{1})\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\cos(\phi^{i}_{2})\\ \vdots\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\ldots\sin(\phi^{i}_{c-3})\cos(\phi^{i}_{c-2})\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\ldots\sin(\phi^{i}_{c-3})\sin(\phi^{i}_{c-2})\end{bmatrix}\,,$ (6)
and take a product-state ansatz for the total wave function
@@
The QdLQA approach is inspired by local quantum annealing bowles_22 , and our goal is to start at the known minimum of one function and simulate a "quantum adiabatic evolution" into the minimum of a cost function for which the solution solves the optimization problem. During the quantum adiabatic evolution, we require that the system remains in a product state. At the end of the simulation, each local state $\vec{\psi}_{i}$ is assigned a color, and Eq. (1) is used to calculate the energy. The scheme is illustrated in Fig. 1.
Figure 1: Three nodes (labeled one, two, and three) connected by three edges. We want to assign three colors (green, red, and blue) so that no two interacting nodes have the same color. The local state is represented by a vector in a three-dimensional sphere (gray arrow). In our initial state, we assign green to node one, and nodes two and three are in the ground states of $-\hat{L}^{x}_{2}$ and $-\hat{L}^{x}_{3}$ respectively (drawn illustratively as a superposition of the three colors). As time $t$ goes from 00 to $1$, Eq. (11) is minimized with gradient descent. During the process, the spins corresponding to node two and three are assigned one eigenstate (color) of $\hat{L}^{z}_{2}$ and $\hat{L}^{z}_{3}$ (corresponding to the highest measurment probability) to provide a solution to Eq. (1).
Figure 2: The $m$ components of the initial distribution of the ground state of $-\hat{L}^{x}_{i}$ in the $\hat{L}^{z}_{i}$ eigenbasis. Each of the components corresponds to a different eigenvalue of $\hat{L}^{z}_{i}$, see Eq. (2), or equivalently, a different color. We show distribution for different numbers of colors $c$ (local Hilbert space dimension).
@@
$E_{\textrm{I}}(\theta)=-\sum\limits_{i\in V\setminus\{j_{\textrm{max}}\}}\expectationvalue*{\hat{L}^{x}_{i}}\,,$ (8)
@@
For our final cost function, we choose the scalar product between probability vectors connected by the edges. This choice arises naturally from our ansatz since two orthogonal $\vec{\psi}_{i}$ correspond to different colors. Using $\vec{p}_{i}$ instead gives the components the interpretation of a measurement probability and emits antiferromagnetic (anti-parallel) solutions:
$E_{\textrm{F}}(\theta)=\sum\limits_{(i,j)\in\mathcal{E}}J_{i,j}\vec{p}_{i}{}^{T}\cdot\vec{p}_{j}\,,$ (9)
where $J_{i,j}=1+h_{i,j}$, and $h_{i,j}$ are drawn from the uniform distribution in the interval $h_{i,j}\in[0,h)$ at each time step during the quantum adiabatic evolution. We note that a similar cost function is used the in graph neural network approaches schuetz_22_2 ; wang_23 . There, however, the $p_{i}$’s emerge from applying the softmax function to the final node embedding. Note that contrary to the algorithm of bowles_22 , this cost function does not have an interpretation as the expectation value of a physical Hamiltonian on the considered quantum product states. Alternatively, we could also have formulated a cost function corresponding to Eq. (1) as a physical Hamiltonian in terms of polynomials in $\hat{L}^{z}_{i}$ and $\hat{L}^{z}_{j}$ deller_23 .
Furthermore, we follow Ref. wang_23 and introduce a term to force the qudits into superpositions of all the colors
$E_{\textrm{W}}(\theta)=\gamma\sum\limits_{i\in V}\vec{p}_{i}{}^{T}\cdot\log(\vec{p}_{i})\,,$ (10)
where $\gamma>0$ is a hyper parameter and the logarithm is taken component wise. For a large $\gamma$. this term favors a finite probability for all possible colors, even those that are initialized close to zero in the ground state of $\expectationvalue*{\hat{L}^{x}_{i}}$ Fig. 2.
@@
$E_{\textrm{Total}}(\theta,t)=(1-t)E_{\textrm{I}}(\theta)+t\big{(}E_{\textrm{F}}(\theta)+E_{\textrm{W}}(\theta)\big{)}\,.$ (11)
In the simulation, we go from $t=0$ to $t=1$ using $N_{steps}$ time steps. At each point in time, we conduct $\textrm{round}(e^{t\alpha})$ gradient descent steps ($\textrm{round}()$ indicates that we round to the nearest integer and $\alpha$ is a hyper parameter). For some of the datasets discussed in Sec. III, we also find it beneficial to not start in the exact ground state of $-L^{x}_{i}$, but rather to start in states where we perturb the initial angles in Eq. (6) to $\phi^{i}_{m}\rightarrow\phi^{i}_{m}+\epsilon_{m}$, where $\epsilon_{m}$ are randomly chosen from the uniform distributed in the interval $[-f,f)$, and $f$ is a hyperparameter. Furthermore, we generate the classical color configuration at each time step and store the one with the lowest value for $E_{\textrm{Potts}}$ [see Eq. (1)], which provides the best configuration found by the algorithm.
For the alternate method focusing entirely on gradient descent, QdGD, each qudit is initialized in a random state with each qudit entry drawn from the uniform distribution in the interval $[0,\tilde{f})$. The state is then normalized, and $E_{\textrm{F}}(\theta)+E_{\textrm{W}}(\theta)$ is minimized directly by gradient descent. In this case, the $h_{i,j}$ are drawn at each gradient descent step and now the total number of steps is given by $N_{steps}$. We also introduce a further hyperparameter, Patience, which stops the algorithm early if no improvement in $E_{\textrm{Potts}}$, Eq. (1), is seen after a certain number gradient descent steps.
@@
To perform the gradient descent, we use the Adam optimizer kingma_14 with learning $\eta$, specifically the PyTorch implementation paszke2019pytorch . All other parameters in the Adam optimizer were set to their PyTorch default values. A detailed list and description of all hyperparameters specific to the algorithms are provided in Appendix A.

### v2

# Qudit-inspired optimization for graph coloring

@@ Abstract
We introduce a quantum-inspired algorithm for graph coloring problems (GCPs) that utilizes qudits in a product state, with each qudit representing a node in the graph and parameterized by $d$-dimensional spherical coordinates. We propose and benchmark two optimization strategies: qudit gradient descent, initiating qudits in random states and employing gradient descent to minimize a cost function, and qudit local quantum annealing, which adapts the local quantum annealing method to optimize an adiabatic transition from a tractable initial function to a problem-specific cost function. Our approaches are benchmarked against established solutions for standard GCPs, showing that our methods not only rival but frequently surpass the performance of recent state-of-the-art algorithms in terms of solution quality and computational efficiency. The adaptability of our algorithm and its high-quality solutions, achieved with minimal computational resources, point to an advancement in the field of quantum-inspired optimization, with potential applications extending to a broad spectrum of optimization problems.
@@ I Introduction
Today, many prominent tasks in industry can be formulated as integer optimization problems, such as how to efficiently use resources kallrath2005solving or do tasks such as job scheduling ku2016mixed or portfolio optimization he2014two . One important example of integer optimization is the graph coloring problem (GCP): given a graph with $V$ nodes and $\mathcal{E}$ edges, assign colors to the nodes such that no two connected nodes share the same color. More generally, one might try to find the minimum number of colors needed for such an assignment to be possible, henceforth referred to as the chromatic number of a graph  lewis_16 . Being NP-complete garey1974some , the GCP can be related to a large number of real-world problems. These include map coloring (e.g., assigning colors to provinces in Spain while avoiding that neighboring provinces have the same color), air traffic control barnier_14 , charging electric vehicles deller_23 , resource allocation (e.g., supply chain or portfolio management), even tasks like designing seating plans, sports tournaments, or sudoku, and managing taxi bookings lewis_16 . Thus, there is much active research into developing algorithms that efficiently find good or optimal solutions to the GCP and scale favorably with the size of the graph.
@@
Quantum technologies might provide one avenue for efficiently solving optimization problems through algorithms such as the Quantum Approximate Optimization Algorithm (QAOA) farhi_14 ; hadfield_19 ; oh_19 or quantum annealing hauke_2020 ; pokharel_21 ; yarkoni_2022 . While there is a straightforward mapping from binary optimization problems to qubit systems lucas_14 ; glover_22 , some integer optimization problems, like GCPs, are more naturally formulated using larger local degrees of freedom. One way to accomplish this in a physics-inspired manner is to extend the notion of a qubit to a $d$-dimensional local degree of freedom, known as a qudit. This, together with the improved capacity for simulations of physical systems with a larger number of local physical degrees of freedom, has inspired the development of hardware implementing qudit systems and qudit-inspired algorithms lu_20 ; wang_20 ; ringbauer_22 . For instance, in Ref. deller_23 , Deller et al. extend QAOA and integer optimization problems, including the GCP, to the qudit regime, and Ref. bottrill_23 explores using qutrits (three-dimensional local degrees of freedom) for the GCP.
@@
Complementary to the improvement of quantum technologies, the connection between quantum physics and optimization problems has enhanced the interest in so-called “physics-inspired algorithms” for optimization. A variety of such algorithms have emerged based on techniques like mean field theory smolin_14 ; veszeli_22 ; bowles_22 , tensor networks bauer_15 ; mugel_22 , neural networks gomes_19 ; Zhao_2021 ; schuetz_22 ; khandoker_2023 , and dynamical evolution Tiunov_19 ; goto_19 ; goto_21 . Some of these new approaches have enjoyed significant success, providing state-of-the-art solutions to existing problems as well as benchmarks to compare against possible quantum-hardware-based solutions.
@@
In this work, we introduce a qudit-inspired algorithm for solving integer optimization problems. First, our approach maps the problem to a qudit system where each qudit is parametrized by $d-$dimensional spherical coordinates. Then, we propose two schemes to solve the optimization problem. In the first scheme, which we refer to as qudit gradient descent (QdGD), each qudit is initialized to a random state and we minimize a cost function via gradient descent. In the second scheme, which we refer to as qudit local quantum annealing (QdLQA), we simulate the quantum annealing from the ground state of a simple Hamiltonian to that of a problem-specific cost function. This is the concept of local quantum annealing introduced by Bowles et al. bowles_22 , in the sense that the system is constrained to remain in a product state through its evolution. We extend this idea to integer optimization, without requiring that the cost function at the end of the evolution has to have an interpretation as a quantum physical Hamiltonian. Our intuition for this approach is that we find the minima of a complicated function $f_{\textrm{final}}(\vec{\phi})$ by first initializing the parameters, $\vec{\phi}$, to a known minimum of a simple function $f_{\textrm{initial}}(\vec{\phi})$. Then, by slowly changing $f_{\textrm{initial}}(\vec{\phi})\rightarrow f_{\textrm{final}}(\vec{\phi})$ while minimizing with respect to the parameters $\vec{\phi}$, an optimal or good solution to $f_{\textrm{final}}(\vec{\phi})$ is found.
@@
We test the algorithms on different standard GCPs and compare to different benchmarks in the literature hebrard_19 ; schuetz_22_2 ; wang_23 . In almost all cases, our approach obtains results as good as or better than recently proposed state-of-the-art algorithms such as graph-neural-network-based approaches schuetz_22_2 ; wang_23 . We also run our algorithms on large sparse datasets where many colors/local degrees of freedom (up to 70 hebrard_19 ) are required to solve the problem. We find that even though our approaches do not find the global minimum for such high-dimensional problems, they still provide relatively good solutions (errors of the order $10^{-3}$) for problem sizes unaccounted for in state-of-the-art works such as schuetz_22_2 ; wang_23 . Notably, in the case of QdLQA, we find that all of these results can be obtained doing only a few gradient descent steps (for most problems only one) at each point in the adiabatic evolution, as was achieved in Ref. bowles_22 .
@@
The rest of this paper is structured as follows. In Sec. II, we introduce the GCP and the algorithms. In Sec. III we benchmark our approach with other state-of-the-art methods. In Sec. IV, we provide a summary and a discussion of improvements and further applications.
@@ II.1 Graph coloring
In our approach, we aim to solve the GCP by reframing it as finding the ground state of the antiferromagnetic Potts model potts_1952 ; wu_82 . The two are closely related wu_88 ; Zdeborova_07 ; schuetz_22_2 , and for an undirected graph $G=(V,\mathcal{E})$ with a set of nodes $V$, edges $\mathcal{E}$, and a set of $c$ colors $\{0,...,c-1\}$, the energy is given by
@@
where each spin is set to one of the colors $\sigma_{i}\in\{0,...,c-1\}$ and $\delta(\sigma_{i},\sigma_{j})=1$ if $\sigma_{i}=\sigma_{j}$ and $\delta(\sigma_{i},\sigma_{j})=0$ otherwise. Thus, this energy is minimized whenever as many spins as possible have a different color to each of their neighbors (so that as many terms as possible are zero). The chromatic number, $\chi(G)$, is defined as the minimum number of colors $c$ required for this cost to reach zero (i.e., when no neighboring vertices share the same color). This problem can be posed in a number of forms, either to simply minimize the cost (NP-complete), to check whether a given graph admits a coloring with a certain number of colors (NP-complete), or to find the chromatic number of a graph (NP-hard) garey1974some .
@@ II.2 Qudit ansatz
|$\ket{\psi}_{i}=\vec{\psi}_{i}=\begin{bmatrix}\cos(\phi^{i}_{0})\\ \sin(\phi^{i}_{0})\cos(\phi^{i}_{1})\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\cos(\phi^{i}_{2})\\ \vdots\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\ldots\sin(\phi^{i}_{c-3})\cos(\phi^{i}_{c-2})\\ \sin(\phi^{i}_{0})\sin(\phi^{i}_{1})\ldots\sin(\phi^{i}_{c-3})\sin(\phi^{i}_{c-2})\end{bmatrix}\,.$ (6)
We choose spherical coordinates for the local states since they automatically keep the state normalized and are easily differentiable, however, other parametrizations are possible (such as using Cartesian coordinates).
Finally, we take a product-state ansatz for the total quantum state
@@
The QdLQA approach is inspired by local quantum annealing bowles_22 , and our goal is to start at the known minimum of one function and simulate a "quantum adiabatic evolution" into the minimum of a cost function for which the solution solves the optimization problem. During the quantum adiabatic evolution, we require that the system remains in a product state. This restriction prevents entanglement buildup in the system, which is believed to lead to the emergence of Barren plateaus McClean_2018 ; Marrero_2021 . At the end of the simulation, each local state $\vec{\psi}_{i}$ is assigned a color, and Eq. (1) is used to calculate the energy. The scheme is illustrated in Fig. 1.
Figure 1: Three nodes (labeled 1, 2, and 3) connected by three edges. We want to assign three colors (green, red, and blue) so that no two interacting nodes have the same color. The local state is represented by a vector in a three-dimensional sphere (gray arrow). In our initial state, we assign green to node 1, and nodes 2 and 3 are in the ground states of $-\hat{L}^{x}_{2}$ and $-\hat{L}^{x}_{3}$ respectively (drawn illustratively as a superposition of the three colors). As time $t$ goes from 00 to $1$, Eq. (11) is minimized with gradient descent. During the process, the spins corresponding to nodes 2 and 3 are assigned one eigenstate (color) of $\hat{L}^{z}_{2}$ and $\hat{L}^{z}_{3}$ (corresponding to the highest measurement probability) to provide a solution to Eq. (1).
Figure 2: The $m$ components of the initial distribution of the ground state of $-\hat{L}^{x}_{i}$ in the $\hat{L}^{z}_{i}$ eigenbasis. Each of the components corresponds to a different eigenvalue of $\hat{L}^{z}_{i}$ [see Eq. (2)], or equivalently, a different color. We show the distribution for different numbers of colors $c$ (local Hilbert space dimension).
@@
$E_{\textrm{I}}(\psi)=-\sum\limits_{i\in V\setminus\{j_{\textrm{max}}\}}\expectationvalue*{\hat{L}^{x}_{i}}\,,$ (8)
@@
For our final cost function, we choose the scalar product between probability vectors connected by the edges. This choice arises naturally from our ansatz since two orthogonal $\vec{\psi}_{i}$ correspond to different colors. Using $\vec{p}_{i}$ instead gives the components the interpretation of a measurement probability and emits antiferromagnetic (antiparallel) solutions:
$E_{\textrm{F}}(\psi)=\sum\limits_{(i,j)\in\mathcal{E}}J_{i,j}\vec{p}_{i}{}^{T}\cdot\vec{p}_{j}\,,$ (9)
where $J_{i,j}=1+h_{i,j}$, and $h_{i,j}$ are drawn from the uniform distribution in the interval $h_{i,j}\in[0,h)$ each time $E_{\textrm{F}}(\psi)$ is called. The reason for introducing the hyperparameter $h$ is to help the simulation escape local minima. If the nodes $i$ and $j$ do not have a color clash, $\vec{p}_{i}{}^{T}\cdot\vec{p}_{j}$ will be approximately 00, and adding $h_{i,j}$ will not matter. However, for $\vec{p}_{i}{}^{T}\cdot\vec{p}_{j}>0$, $h_{i,j}$ will increase the interaction strength between the nodes and encourage a change of colors. Furthermore, we note that a similar cost function (without the $h$ hyperparameter) is used the in graph neural network approaches schuetz_22_2 ; wang_23 . There, however, the $\vec{p}_{i}$ emerge from applying the softmax function to the final node embedding. Note that contrary to the algorithm of bowles_22 , this cost function does not have an interpretation as the expectation value of a physical Hamiltonian on the quantum product states considered. Alternatively, we could also have formulated a cost function corresponding to Eq. (1) as a physical Hamiltonian in terms of polynomials in $\hat{L}^{z}_{i}$ and $\hat{L}^{z}_{j}$ deller_23 .
Furthermore, we follow Ref. wang_23 and introduce a term to force the qudits into superpositions of all the colors,
$E_{\textrm{W}}(\psi)=\gamma\sum\limits_{i\in V}\vec{p}_{i}{}^{T}\cdot\log(\vec{p}_{i})\,,$ (10)
where $\gamma>0$ is a hyperparameter and the logarithm is taken component wise. For large $\gamma$, this term favors a finite probability for all possible colors, even those that are initialized close to zero in the ground state of $-\hat{L}^{x}_{i}$ (see Fig. 2). Heuristically, we find that this term helps to avoid getting stuck in local minima (similar to what was observed with graph neural networks in Ref. wang_23 ).
@@
$E_{\textrm{Total}}(\psi,t)=(1-t)E_{\textrm{I}}(\psi)+t\big{(}E_{\textrm{F}}(\psi)+E_{\textrm{W}}(\psi)\big{)}\,.$ (11)
In the simulation, we go from $t=0$ to $t=1$ using $N_{\text{steps}}$ time steps, and at each point in time, we conduct $\alpha$ gradient descent steps. For some of the datasets discussed in Sec. III, we also find it beneficial not to start in the exact ground state of $-L^{x}_{i}$, but rather to start in states where we perturb the initial angles in Eq. (6) to $\phi^{i}_{m}\rightarrow\phi^{i}_{m}+\epsilon_{m}$, where $\epsilon_{m}$ are randomly chosen from the uniform distributed in the interval $[-f,f)$, and $f$ is a hyperparameter. Furthermore, we generate the classical color configuration at each time step and store the one with the lowest value for $E_{\textrm{Potts}}$ [see Eq. (1)], which provides the best configuration found by the algorithm.
For the alternative method focusing entirely on gradient descent, QdGD, each qudit is initialized in a random state with each qudit entry drawn from the uniform distribution in the interval $[0,\tilde{f})$. The state is then normalized, and $E_{\textrm{F}}({\psi})+E_{\textrm{W}}({\psi})$ is minimized directly by $N_{\text{steps}}$ gradient descent steps. We also introduce a further hyperparameter, Patience, which stops the algorithm early if no improvement in $E_{\textrm{Potts}}$ [Eq. (1)] is seen after a certain number gradient descent steps.
@@
To perform the gradient descent, we use the Adam optimizer kingma_14 with learning $\eta$, specifically the PyTorch implementation paszke2019pytorch . All other parameters in the Adam optimizer were set to their PyTorch default values. Also, note that we provide pseudocode for the two algorithms in Appendix A, an overview of runtimes for the calculations presented later in Appendix B, and a detailed list and description of all hyperparameters specific to the algorithms in Appendix C.